What’s the mole fraction of oxygen in a solution containing 11. mg/L oxygen (as O2) and 150. mg/L of NaCl? Solve the problem exactly, assuming the NaCl dissolves, and then using the approximation that all the ion molar concentrations are negligibly small compared to water.

ANSWER: 6.2e-6

JUST SHOW STEPS PLEASE AND THANKS :-)
Answer #1

volume of water = 1 L = 1000 ml

mass of water = 1000 g as density of water = 1.0 g / ml

mole of water = 1000 g / 18 g / mole = 55.5556 mole.

mole of O2 = 0.011 g / 32 g / mole = 3.475 * 10^-4 mole.

and

mole of NaCl = 0.150 g / 58.5 g / mole = 2.562 * 10^-3 mole.

total mole = (55.5556 + 3.475 * 10^-4 + 2.562 * 10^-3) = 55.55846 mole.

mole fraction of O2 = 3.475 * 10^-4 / 55.55846 = 6.2 * 10^-6 (answer)